# Safe Explicable Policy Search

### v1

@@ VI-A Safety
We conduct this evaluation on the safety-gym’s PointGoal task as shown in Fig. 4(a). Here, the agent is required to reach the goal (green) while avoiding stepping into hazardous areas (dark blue) and hitting vases (light blue). In the agent’s model, the task objective is to reach the goal efficiently. Furthermore, there are constraints on the costs associated with hazards and vases. We assume the user’s belief of the agent’s objective includes avoiding moving close to vases as they are considered fragile. However, the user is unaware of the safety risk associated with the agent entering the hazardous areas.
We evaluated SEPS against the agent’s performance in its own model ($\mathcal{M}_{A}$), under user expectations ($u_{H}$), and under EPS ($EPS$). Here, no safety constraints are incorporated into the baselines $u_{H}$ and $EPS$, to emphasize the need for safety constraints in explicable planning. In the next evaluation, we will incorporate the consideration of safety constraint into these baselines to compare with SEPS. We compared the behaviors generated and the returns obtained under the surrogate reward, the agent’s task reward, and the cost function.
Results and Discussion
The results of this evaluation are shown in Fig. 2. The main result is in the violation of safety constraints reflected in the returns under the cost function $\mathcal{C}_{1}$ shown in Fig. 2(c). The closer the cost return to 00, the safer the behavior is. $u_{H}$ and EPS resulted in more safety violations than SEPS and the optimal behavior (which happened to incur a low cost due to the environment setting). This is due to the fact that the surrogate objective only includes consideration of the vases. EPS incurred the most cost since it sometimes resulted in the agent wandering around without reaching the goal, leading to the low return in Fig. 3(b) and demonstrating the instability issue due to linearly combining objectives. Even though the agent in SEPS executed a curvy path to the goal, its return in $\mathcal{R}_{A}$ is almost as high as the optimal behavior (Fig. 3(b)). This is because the task objective in the agent’s model features a penalty proportional to its displacement from the goal. As long as the agent makes progress towards the goal (whether straightly or in a curve), little penalty would be given.
(a) PointGoal
(b) PointButton
Figure 4: Lines show resulting trajectories of different methods in the two simulation domains; SEPS-blue, $\mathcal{M}_{A}$/$\mathcal{R}_{A}{-}CPO$-red, $u_{H}$/$u_{H}{-}CPO$-orange, and $EPS$/$EPS{-}CPO$-green.
Fig. 4(a) shows the behaviors generated in the PointGoal task. As user model is assumed to not associate risk with hazardous regions but only with vases, $u_{H}$ and $EPS$ generated unsafe trajectories that passed through hazardous regions. In $\mathcal{M}_{A}$, the agent followed the most efficient path to reach the goal without colliding with the vases or entering the hazardous regions, although such a path narrowly passing through the two vases may be seen as risky. $SEPS$ also avoided hazardous regions and staying further away from the vases, to be explicable.
VI-B Efficiency
In our second evaluation, our aim is to answer ‘Is the agent efficient under its own task objective?’ and ‘Is SEPS’s policy explicable and safe while being efficient?’. We designed an evaluation scenario where the user’s model aligned with the safety requirements of the agent but differed from the agent’s task objective.
We conducted this evaluation in a modified version of the safety-gym’s PointButton task as shown in Fig. 4(b). Here, the agent is required to reach the goal (green) while avoiding contact with moving gremlins (purple). In addition, there are two buttons (orange) in the environment that can be pressed for a small reward. In the agent’s model, the task objective is to reach the goal efficiently subject to constraints on the costs associated with gremlins. We assumed that the user’s belief of the agent’s objective was to press all the buttons to maximize its reward before reaching the goal while avoiding the gremlins.
We evaluated SEPS against three baselines that enforce safety constraints via CPO: the agent’s objective $\mathcal{R}_{A}-CPO$, the user’s expectations (surrogate reward) $u_{H}-CPO$, and Explicable Policy Search $EPS-CPO$. We compared the behaviors generated and the returns obtained under the surrogate reward, the agent’s task reward, and the cost function.
Results and Discussion
The results for this evaluation are shown in Fig. 3. As $u_{H}-CPO$ solely optimizes the surrogate objective ($u_{H}$), it performed poorly under the agent’s task objective ($\mathcal{R}_{A}$). Similarly, as $\mathcal{R}_{A}-CPO$ solely optimizes $\mathcal{R}_{A}$, it performed poorly under $u_{H}$. Explicable Policy Search (EPS), when optimized subject to safety constraints ($EPS-CPO$), achieved better performance under the surrogate reward but performed poorly under the agent’s task objective due to the lack of any constraint on the objective. In $EPS-CPO$ we set the reconciliation parameter $\lambda=3$. Empericially, we found that $\lambda=3$ balanced reasonably well between the surrogate reward and the agent’s reward.
(a) $\tau_{\mathcal{M}_{A}}$
(b) $\tau_{\mathcal{M}_{A}^{H}}$
@@ VI-B Efficiency
In $SEPS$, we set $d_{0}=0.0$ and $d_{1}=2.5$. Note that these bounds have clear semantic meanings (e.g., the suboptimality level under the respective metric), which makes them easier to specify than the hyper-parameters in EPS. It is worth noting that unlike $SEPS$, all baselines have a single constraint on safety. Addressing multiple constraints is more challenging since they may sometimes conflict with each other. Even under such a case, $SEPS$ stayed close to $\mathcal{R}_{A}-CPO$ under both $\mathcal{R}_{A}$ and the safety cost. In other words, it is able to remain explicable while sufficiently satisfying both the performance and safety requirements.
The behaviors generated by different methods for an example seed in the PointButton task are shown in Fig. 4(b). In SEPS and $EPS-CPO$ the agent pressed a single button before reaching the goal. However, in $EPS-CPO$ the agent showed a suboptimal behavior by taking a longer path to the button and going around the button before reaching the goal. As expected, in $u_{H}-CPO$ the agent pressed both buttons while in $\mathcal{R}_{A}-CPO$ it headed straight to the goal.
VI-C Physical Robot Experiment
In this experiment, our aim is to analyze the behaviors generated when a physical robot (UR5) assists a user to set up a dining table. We designed the experiment (see Fig. 5) in a discrete setting where the location of the objects was preset to specific waypoints, and the robot’s movement between them was calibrated using the UR5’s inverse kinematics library.
We assume that the user’s expectations differ from the robot’s functionality. The robot is required to pass the cups to the user by placing them on the dining table. The user expects the robot to place the cups close to where he is standing. However, doing so may tip the glass in front of the robot. Tipping the glass will spill its contents, which is unsafe in the robot’s model and has a high cost. We analyzed the behavior generated by SEPS and compared it with the optimal behavior in its own model, the user’s expected behavior, and $EPS{-}CPO$.
(a) $\tau_{EPS}$
(b) $\tau_{SEPS}$
Figure 6: Safe explicable behaviors generated by $EPS{-}CPO$ and SEPS.
Results and Discussion
The robot picked the cup and placed it on the table next to the glass when executing the optimal behavior in its model (see Fig. 5(a)). When executing the user’s expected behavior, the robot tipped the glass (replaced by an empty plastic bottle in this execution for safety reasons), as shown in Fig. 5(b). In reality, the robot would have spilled the liquid as the glass obstructed the manipulator. With EPS (see Fig. 6(a)), the robot first picked up the glass and moved it across the table. Then, it picked up the cup and placed it close to the user to meet his expectation. However, picking and placing the glass was less efficient due to additional movements. With SEPS (see Fig. 6(b)), the robot first moved the manipulator to push the glass gently across the table to clear it out of the way. Then, it picked up the cup and placed it close to the user as in EPS. While both EPS and SEPS generated explicable behaviors, SEPS’s is more efficient, thanks to the constraint on the task objective.
@@ VII Conclusion
In this paper, we introduced the problem of Safe Explicable Policy Search (SEPS), which aimed to search for explicable policies while considering both safety and task efficiency. We adopted and adapted the explicability measure in EPS to formulate SEPS as a CMDP. SEPS relaxed the need for specifying multiple hyper-parameters for explicable policy search by considering the agent’s task objective as a separate constraint. These treatments allowed us to apply CPO to solve SEPS to satisfy the safety and quality requirements during the learning process while maximizing explicability. We derived an analytical solution for the most common case of SEPS with two constraints. We evaluated SEPS for safety and efficiency in safety-gym environments. The results demonstrated that SEPS was effective in learning safe, explicable, and efficient policies. We also demonstrated the applicability of our approach via a physical robot experiment.

### v3

@@ VI-B Ablation Study
In $\mathcal{M}_{A}$, the agent must reach the goal (green) without colliding with gremlins (purple), which move in fixed patterns, and may optionally press buttons (orange). The reward $\mathcal{R}_{A}$ assigns positive values for reaching the goal and pressing buttons and negative values for moving away from the goal, while the cost $\mathcal{C}_{1}$ penalizes collisions with gremlins. According to user expectations ($u_{H}$), the agent must press both buttons before reaching the goal. Thus, $u_{H}$ mirrors $\mathcal{R}_{A}$ except that deviations from the goal do not incur a negative value.
[…]
In SEPS, we set $d_{0}=0.0$ and $d_{1}=2.5$. These limits have intuitive semantic interpretations. For example, the agent can deviate from the goal provided that its expected return in $\mathcal{R}_{A}$ remains above $0.0$, making them easier to specify than the hyperparameters in EPS. Unlike SEPS, all baselines impose only a single safety constraint, whereas handling multiple constraints can be more challenging due to potential conflicts. Even under such conditions, SEPS remained close to $\mathcal{M}_{A}$ in both $\mathcal{R}_{A}$ and $\mathcal{C}_{1}$, demonstrating that it can remain explicable while meeting both performance and safety requirements.
VI-C Physical Robot Experiment
In this experiment, we analyzed the behaviors generated when a physical UR5 robot assisted a user to set a dining table, to demonstrate that SEPS also applies to task planning domains. The experiment (Fig. 5) was designed in a discrete setting with preset object locations, and robot motions calibrated using the UR5 inverse kinematics library. The states were defined by the locations of objects (A, B, C, D, E, shown in Fig. 5(a)) and their status (steady or slipped).
(a) $\tau_{\text{AGT}}$
(b) $\tau_{\text{HUM}}$
@@ VI-C Physical Robot Experiment
[…]
(a) $\tau_{\text{SEPS}+\mathcal{R}_{A}{-}C_{0}}$
(b) $\tau_{\text{SEPS}}$
[…]
@@ VII Conclusion
In this paper, we introduced Safe Explicable Policy Search (SEPS), which addresses the problem of searching for explicable policies when both the environment model and user expectations are unknown, while ensuring safety and efficiency. We adopted the surrogate reward model from EPS as the explicability metric in SEPS, which allowed a reduction of optimization in two models to a single model. This reduction allowed us to formulate the problem as a CMDP, which simplifies the specification of safety: the thresholds for the expected returns in CMDP are relatively straightforward to set as opposed to tuning weights and sensitivity parameters in a risk-aware RL framework. Similarly, unlike EPS, SEPS alleviates the need for multiple hyperparameters by treating the agent’s task objective as a constraint. The CPO method elegantly solves CMDP problems, owing to its convergence and safety guarantees, but is limited to a single constraint. We addressed this by deriving an analytical solution to SEPS for the common case with two constraints. Our evaluation in Safety-Gym environments showed that SEPS learned safe, explicable, and efficient policies, and we further demonstrated its applicability through a physical robot experiment.
Meanwhile, there are several limitations with SEPS to consider in future work. The complexity of the solution increases with the number of constraints, which remains to be studied when there are more than two constraints. Imposing both the suboptimality and safety constraints can lead to search failures. In such cases, feasible solutions can be obtained by relaxing the suboptimality criterion. How to bound such relaxations is a challenging problem. Since CPO ensures safety only in expectation with respect to state visitations, agents may still violate safety constraints during and after learning with stricter safety criteria, such as trajectory-level constraints. Addressing this limitation, for instance, through shielding methods [44], remains an interesting direction for future work.